OpenStudy (anonymous):

Give an example for a sequence \(A_{n}\) of closed sets \(A_{n} \subset \mathbb{R}\) such that the union \(\bigcup_{n=0}^{\infty} A_{n}\)is not closed. Give an example for a sequence \(U_{n}\)of open sets \(U_{n} \subset \mathbb{R}\), such that the intersection \(\bigcap_{n=0}^{\infty} U_{n}\)is not open

OpenStudy (anonymous):

\[ A_n=\{ \frac 1 n\} \] is closed

OpenStudy (anonymous):

\[ \cup A_n \] is not closed. Why?

OpenStudy (anonymous):

\[ U_n =(-\frac 1 n , \frac 1 n) \] is open.

OpenStudy (anonymous):

ok..

OpenStudy (anonymous):

\[ \cap U_n =\{0\} \] is not open.

OpenStudy (anonymous):

hmm

OpenStudy (anonymous):

can it be a solution for our question or is it like a hint what u wrote there?

OpenStudy (anonymous):

No, it is the solution.

OpenStudy (anonymous):

cool so short, thank you very much... i hope i can understand it soon by myself how it done..

OpenStudy (anonymous):

yw